### v1

# Self is the Best Learner: CT-free Ultra-Low-Dose PET Organ Segmentation via Collaborating Denoising and Segmentation Learning

Abstract
Organ segmentation in Positron Emission Tomography (PET) plays a vital role in cancer quantification. Low-dose PET (LDPET) provides a safer alternative by reducing radiation exposure. However, the inherent noise and blurred boundaries make organ segmentation more challenging. Additionally, existing PET organ segmentation methods rely on co-registered Computed Tomography (CT) annotations, overlooking the problem of modality mismatch. In this study, we propose LDOS, a novel CT-free ultra-LDPET organ segmentation pipeline. Inspired by Masked Autoencoders (MAE), we reinterpret LDPET as a naturally masked version of Full-Dose PET (FDPET). LDOS adopts a simple yet effective architecture: a shared encoder extracts generalized features, while task-specific decoders independently refine outputs for denoising and segmentation. By integrating CT-derived organ annotations into the denoising process, LDOS improves anatomical boundary recognition and alleviates the PET/CT misalignments. Experiments demonstrate that LDOS achieves state-of-the-art performance with mean Dice scores of 73.11% (18F-FDG) and 73.97% (68Ga-FAPI) across 18 organs in 5% dose PET. Our code is publicly available.
Keywords:
Organs segmentation Ultra-LDPET MAE Denoising.
1 Introduction
Positron emission tomography (PET) is a powerful molecular imaging that visualizes radiotracer distribution to reveal physiological processes. Low-dose PET (LDPET) reduces radiation exposure and has demonstrated diagnostic equivalence to Full-Dose PET (FDPET) [25, 26, 14, 4]. However, its clinical adoption remains limited by the lack of robust quantitative tools in high-noise settings. Accurate tracer uptake measurements in non-tumoral organs are critical for tumor quantification within Volumes of Interest (VOIs) [1, 15]. LDPET organ segmentation offers a promising approach for tracer uptake assessment and kinetic measurements [23, 27], yet it remains relatively understudied.
Figure 1: (a) Misalignments between PET and CT caused by respiratory motion (e.g., liver displacement). (b) Noise and blurred boundary artifacts in LDPET.
In this study, we summarize the challenges of LDPET organ segmentation: (1) Data scarcity: Clinical FDPET dominance limits LDPET dataset availability. (2) Annotation complexity: Low soft-tissue contrast in PET, worsened in LDPET, makes organ annotation challenging. Existing PET organ segmentation methods rely on co-registered Computed Tomography (CT) annotations, overlooking the issue of modality mismatch (Fig. 1a). (3) Noise-induced ambiguity: Blurred anatomical boundaries hinder accurate segmentation (Fig. 1b).
To address these challenges, we propose LDOS, a novel CT-free LDPET organ segmentation pipeline via collaborating denoising and segmentation learning. Inspired by MAE [8], we reinterpret LDPET as a naturally masked version of FDPET, where the denoising process inherently recovers organ-level semantics. Recent studies have demonstrated that incorporating organ annotations into LDPET denoising can improve FDPET reconstruction, highlighting the shared semantic priors between denoising and segmentation tasks [11, 24, 5].
Unlike traditional MAE-based methods that depend on large-scale pretraining, LDOS employs self-denoising to reinforce organ-level semantic learning, reducing the risks of overfitting and bias-learning in small-scale PET datasets [21, 16, 18]. From a technical perspective, LDOS uses a simple yet effective training pipeline. A shared encoder extracts generalized features for both tasks, while task-specific decoders independently refine outputs for denoising (supervised by FDPET) and segmentation (supervised by CT annotations during training only). By incorporating CT-derived organ annotations into the denoising process, LDOS improves anatomical boundary recognition and mitigates PET/CT misalignments.
Figure 2: Overview of LDOS. LDOS employs the shared encoder and independent decoders for denoising and segmentation tasks.
2 Related Work
2.0.1 PET Organ Segmentation
Existing PET segmentation methods primarily focus on organs with high tracer uptake (e.g., tumors) in FDPET, while organ segmentation in LDPET remains largely unexplored [20, 17, 2]. Furthermore, PET organ segmentation methods rely on co-registered CT annotations, ignoring modality mismatch [19, 22]. To address this gap, we propose LDOS, which operates directly on noisy ultra-LDPET data. By incorporating CT-derived organ annotations into the denoising process, LDOS improve anatomical boundary recognition and mitigates PET/CT misalignments.
2.0.2 MAE in Medical Image Analysis
MAE excels in natural images with large-scale pretraining (e.g., ImageNet [6]). However, medical imaging lacks comparable datasets, making pretraining on limited data prone to overfitting and bias [21, 16, 18]. Inspired by MAE, where masking removes random patches, we reinterpret LDPET as positional masking of FDPET, where low-dose acquisition inherently “masks” high-frequency details. Instead of resource-intensive pretraining, LDOS integrates denoising and segmentation into a single-stage collaborative learning process, avoiding small-data pitfalls and ensuring anatomical consistency.
3 Main Methodology
LDOS employs a simple yet effective training pipeline, which improves segmentation accuracy by jointly reconstructing FDPET signals. As shown in Fig. 2, it employs a shared encoder $E_{\theta}$ and two task-specific decoders $D_{\phi}$ and $D_{\psi}$.
3.1 Model Architecture
3.1.1 Shared Encoder
LDOS uses the nnU-Net backbone [9] with the ResEncL configuration [10]. Inspired by MAE, LDOS employs the LDPET denoising process to learn semantic latent representations. MAE uses the patch-level discretized images as inputs [7]. In the medical imaging context, large-scale patch masking leads to excessive information loss. In our work, LDOS is regarded as a pixel-level masking approach to retain finer details:
${{x}_{i}}=\psi(\Lambda{{r}_{i}},k)$ (1)
[…]
$F({{x}_{i}})=E({{\vartheta}_{e}};{{x}_{i}})\odot\zeta({{x}_{i}})$ (2)
where $F({{x}_{i}})$ represents shared semantic features, $E$ is the encoder, ${{\vartheta}_{e}}$ denotes encoder parameters, $\odot$ signifies dimensionality stacking, and $\zeta$ is the residual connection. Unlike standard MAE approaches that focus on visible pixels, our method models all pixels, providing comprehensive semantic representation.
3.1.2 Segmentation and Denoising Decoders
The encoder output feeds into both segmentation and denoising decoders, which share a consistent architecture. To enhance feature learning, a deep supervision strategy is incorporated:
$P_{full-dose}^{z}=D(\vartheta_{d}^{z};F({{x}_{i}}))$ (3)
$P_{seg}^{z}=\frac{\exp({{D}_{c}}(\vartheta_{s}^{z};F({{x}_{i}})))}{\sum\limits_{C}{\exp({{D}_{c}}(\vartheta_{s}^{z};F({{x}_{i}}))}}$ (4)
where $P_{full-dose}^{z}$ is the FDPET prediction at scale $z$, $D$ is the decoder, and $\vartheta_{d}^{z}$ denotes the decoder parameters. Similarly, $P_{seg}^{z}$ refers to the segmentation prediction for the $c$-th class, with ${{D}_{c}}$ being the decoder for class $c$ and $C$ the total number of segmentation classes.
3.2 Training via Collaborative Denoising and Segmentation
To address PET/CT misalignments [12, 2], misalignment data augmentation [19] is applied during training. This strategy mitigates label inconsistencies and improves coordination between segmentation and denoising. The denoising decoder is supervised by the loss function ${{L}_{g}}$:
${{L}_{g}}=\sum\limits_{Z}{{{\omega}^{z}}\left|P_{full-dose}^{z}-r\right|}$ (5)
where $Z$ is the total number of scales, ${{\omega}^{z}}$ is the weight of scale $z$, and $r$ is the FDPET image. For segmentation, a combination of Dice and cross-entropy losses is used:
${{L}_{ce}}=\sum\limits_{Z}{\sum\limits_{C}{{{\upsilon}^{z}}\left(-{{y}_{c}}\log{{P}_{seg,c}^{z}}-(1-{{y}_{c}})\log(1-{{P}_{seg,c}^{z}})\right)}}$ (6)
${{L}_{dice}}=\sum\limits_{Z}{{{\upsilon}^{z}}\left(1-\frac{2\sum\limits_{C}{({{y}_{c}}\cdot\rho(soft\max({{P}_{seg,c}^{z}})))}}{\sum\limits_{C}{{{y}_{c}}+\rho(soft\max({{P}_{seg,c}^{z}}))}}\right)}$ (7)
${{L}_{s}}={{L}_{ce}}+{{L}_{dice}}$ (8)
where ${{L}_{ce}}$ is the cross-entropy loss, $C$ is the total number of classes, ${{\upsilon}^{z}}$ is the weight at scale $z$, ${{y}_{c}}$ is the ground truth for class $c$, $\rho$ is one-hot embedding, and ${{P}_{seg,c}^{z}}$ is the segmentation prediction. The total loss function combines denoising and segmentation losses:
${{L}_{total}}=\varphi{{L}_{g}}+\lambda{{L}_{s}}$ (9)
where $\varphi=e^{-\ell E}$; $\ell=0.002$, $\lambda=1$, and $E$ denotes the number of epochs. Initially, denoising is emphasized to capture semantic features, and as training progresses, the focus shifts to refine segmentation accuracy.
Table 1: Performance of the 18 organs segmentation results on the 18F-FDG and 68Ga-FAPI datasets (5-fold cross validation). Dice, IoU, HD95, and ASD are used to evaluate LDOS, and ‘±’ represents ‘mean ± standard deviation’.The green and yellow highlights mean the highest and lowest segmentation results.
Organs 18F-FDG Dataset 68Ga-FAPI Dataset
Dice (%) IoU (%) HD95 (mm) ASD (mm) Dice (%) IoU (%) HD95 (mm) ASD (mm)
Spleen 76.28 ± 2.31 58.99 ± 3.50 13.56 ± 4.32 3.77 ± 1.16 69.37 ± 5.61 54.90 ± 5.85 17.67 ± 4.32 4.95 ± 0.88
Colon 62.22 ± 3.47 53.33 ± 4.94 32.35 ± 8.07 9.42 ± 2.54 44.76 ± 4.79 29.68 ± 1.61 42.99 ± 7.33 8.58 ± 2.14
Urinary Bladder 72.75 ± 7.64 58.61 ± 11.99 9.23 ± 5.87 3.65 ± 3.16 79.94 ± 2.99 66.15 ± 6.51 14.74 ± 8.06 3.50 ± 1.44
Sacrum 69.20 ± 4.18 57.84 ± 5.03 12.29 ± 2.72 3.92 ± 0.65 71.05 ± 4.97 53.55 ± 3.86 10.20 ± 2.56 2.92 ± 0.21
Vertebrae 77.38 ± 4.85 65.87 ± 2.19 10.06 ± 3.31 2.84 ± 0.96 72.10 ± 2.22 54.92 ± 0.78 11.06 ± 6.49 2.40 ± 0.13
Heart 81.06 ± 4.30 72.59 ± 5.35 12.11 ± 3.56 4.10 ± 2.07 82.51 ± 1.57 69.92 ± 1.59 10.97 ± 0.86 3.71 ± 0.46
Aorta 76.38 ± 3.18 59.63 ± 2.04 12.20 ± 2.84 4.34 ± 0.75 72.32 ± 4.66 58.12 ± 2.10 10.44 ± 1.85 2.91 ± 0.20
Clavicula 63.34 ± 5.09 54.53 ± 4.38 15.78 ± 7.34 5.57 ± 2.76 59.08 ± 5.49 40.95 ± 3.75 10.39 ± 3.09 2.85 ± 0.10
Femur 77.05 ± 2.46 66.71 ± 3.94 10.88 ± 4.34 3.38 ± 1.94 88.58 ± 1.67 79.91 ± 2.05 5.74 ± 1.40 2.53 ± 0.95
Hip 76.50 ± 4.00 65.70 ± 3.78 10.54 ± 6.84 3.77 ± 1.64 76.86 ± 2.08 62.75 ± 2.60 8.67 ± 2.00 2.70 ± 0.19
Autochthon 75.92 ± 2.06 63.69 ± 2.75 8.28 ± 4.34 3.16 ± 1.55 85.05 ± 3.70 74.08 ± 3.22 5.56 ± 0.61 2.07 ± 0.12
Kidney 75.61 ± 4.79 64.35 ± 3.97 14.35 ± 5.06 4.78 ± 1.62 75.81 ± 3.62 62.07 ± 3.47 11.03 ± 0.58 3.30 ± 0.28
Brain 94.63 ± 3.65 91.12 ± 2.39 4.58 ± 1.79 1.87 ± 0.68 95.64 ± 0.87 91.65 ± 0.70 4.57 ± 0.23 1.86 ± 0.15
Liver 86.52 ± 2.46 78.78 ± 3.57 14.82 ± 5.33 5.14 ± 2.25 85.19 ± 1.97 72.30 ± 3.48 19.80 ± 5.50 5.56 ± 1.38
Stomach 63.90 ± 8.16 52.76 ± 7.49 23.24 ± 12.56 7.59 ± 2.01 69.76 ± 8.80 57.94 ± 8.86 29.92 ± 11.75 7.50 ± 2.44
Pancreas 52.61 ± 2.33 40.64 ± 5.94 18.80 ± 7.66 5.45 ± 2.29 54.58 ± 5.50 30.58 ± 6.94 21.38 ± 1.94 4.67 ± 0.42
Lung 89.33 ± 0.80 84.72 ± 2.01 12.10 ± 2.48 4.21 ± 1.50 91.88 ± 1.74 84.55 ± 2.03 7.81 ± 0.70 2.57 ± 0.25
Esophagus 50.78 ± 2.14 37.89 ± 6.45 19.49 ± 9.56 7.32 ± 5.32 56.96 ± 2.44 41.51 ± 4.13 9.00 ± 1.74 2.33 ± 0.17
Mean 73.11 ± 1.02 62.65 ± 3.07 14.15 ± 2.04 4.68 ± 1.01 73.97 ± 2.26 60.31 ± 2.72 14.00 ± 2.17 3.72 ± 0.46
4 Experiment and Results
4.1 Datasets and Implementation Details
LDOS was evaluated on datasets of images from patients injected with 18F-FDG and 68Ga-FAPI tracers, with sampling times of 300 seconds (100% dose) for FDPET and 15 seconds (5% dose) for ultra-LDPET. Each dataset included 18 segmented organs and was split into training (80%), validation (10%), and testing (10%) subsets.
LDOS was implemented based on the nnU-Net and trained from scratch. The patch size was $192\times 192\times 192$, and the batch size was 2. Training was conducted on an NVIDIA GeForce RTX 4090 GPU with 24 GB memory.
Table 2: Performance of the 18 organs segmentation results on real LDPET and FDPET images. Dice (%) is used as the metric.
Metric 18F-FDG 68Ga-FAPI
FDPET LDPET LDOS FDPET LDPET LDOS
Dice 76.03 ± 3.61 68.03 ± 2.35 73.11 ± 1.02 74.65 ± 3.09 65.73 ± 2.79 73.97 ± 2.26
IoU 62.98 ± 4.01 50.77 ± 7.39 62.65 ± 3.07 63.68 ± 2.45 48.70 ± 8.31 60.31 ± 2.72
HD95 15.71 ± 3.50 21.05 ± 3.84 14.15 ± 2.04 11.03 ± 2.64 23.94 ± 5.33 14.00 ± 2.17
ASD 3.96 ± 1.67 6.88 ± 1.78 4.68 ± 1.01 3.12 ± 1.06 7.06 ± 2.00 3.72 ± 0.46
Figure 3: A sample of head and abdominal organs segmentation on the 18F-FDG. For clarity, segmentation results for some organs are not displayed.
4.2 Multiorgan Segmentation Results
The segmentation performance of LDOS was evaluated using multiple metrics, including Dice Similarity Coefficient (Dice), Intersection over Union (IoU), 95th percentile Hausdorff distance (HD95), and Average Surface Distance (ASD). A 5-fold cross-validation was conducted on both datasets. The quantitative segmentation results are shown in Table 1. The average Dice scores across the 18 organs were $73.11\%\pm 1.02\%$ and $73.97\%\pm 2.26\%$ for the 18F-FDG and 68Ga-FAPI datasets, respectively. Additionally, the average HD95 were $14.15\,\text{mm}\pm 2.04\,\text{mm}$ and $14.00\,\text{mm}\pm 2.17\,\text{mm}$, demonstrating the robust segmentation capability of LDOS at ultra-LPPET (5% dose).
LDOS was compared to the segmentation results of real LDPET and FDPET. For a fair comparison, the same nnU-Net architecture was employed, excluding the denoising decoder loss. As shown in  2 and Fig. 4, LDOS achieved notable improvements in Dice scores, with $5.08\%$ and $8.24\%$ higher accuracy than LDPET on the 18F-FDG and 68Ga-FAPI datasets, respectively. Moreover, LDOS achieved segmentation accuracy comparable to or surpassing FDPET for certain organs.
Figure 4: Performance comparison of selected organs across different methods on the 18F-FDG (a) and 68Ga-FAPI (b) datasets. Dice (%) is used as the metric.
Table 3: Performance comparison of different methods on the 18F-FDG and 68Ga-FAPI datasets. Dice (%) is used as the metric. The best performance is highlighted in bold, and the suboptimal performance is underlined.
Method Pre-train Data 18F-FDG 68Ga-FAPI
Baseline (ResEncU) - $68.03\pm 2.35$ $65.73\pm 2.79$
EM-Net [3] - $65.45\pm 5.11$ $62.77\pm 3.46$
nnU-Net [9] - $67.58\pm 7.30$ $63.96\pm 4.15$
AutopetIII $68.86\pm 3.61$ $64.86\pm 1.98$
S3D (nnU-Net) [21] Our Datasets $70.35\pm 1.78$ $70.87\pm 2.35$
AutopetIII and Our Datasets $70.03\pm 0.96$ $72.76\pm 2.36$
- $67.39\pm 5.00$ $64.97\pm 6.27$
AMAES (U-Net B) [16] BRAINS-45K $66.92\pm 3.76$ $64.68\pm 2.99$
AutopetIII $68.45\pm 4.37$ $64.56\pm 3.89$
AutopetIII and Our Datasets $71.01\pm 1.65$ $69.33\pm 1.24$
LDOS - $73.11\pm 1.02$ $73.97\pm 2.26$
Figure 5: Visualization of segmentation results on the 18F-FDG dataset.
PET organ annotations for model training are derived from corresponding CT images, as direct organ labeling on PET images is impractical. However, achieving precise alignment of training is challenging due to inherent spatial shifts. As illustrated in Fig. 3, LDOS alleviates the misalignments by semantic feature learning of the self-denoising process, providing a novel solution to improve ultra-LDPET organ segmentation.
Table 4: Ablation studies of LDOS. LDPET-LDOS/ow means LDOS without the denoising process, and Fake-FDPET means that denoising and segmentation are performed sequentially.The best performance is highlighted in bold.
Method Dataset Dice (%) IoU (%) HD95 (mm) ASD (mm)
LDPET 18F-FDG $68.03\pm 2.35$ $50.77\pm 7.39$ $21.05\pm 3.84$ $6.88\pm 1.78$
Fake-FDPET - $73.75\pm 3.02$ $61.46\pm 2.96$ $13.76\pm 3.05$ $5.35\pm 1.84$
LDPET-LDOS/ow - $72.01\pm 4.00$ $61.01\pm 2.66$ $15.02\pm 2.59$ $5.72\pm 1.55$
LDPET-LDOS - $73.11\pm 1.02$ $62.65\pm 3.07$ $14.15\pm 2.04$ $4.68\pm 1.01$
LDPET 68Ga-FAPI $65.73\pm 2.79$ $48.70\pm 8.31$ $23.94\pm 5.33$ $7.06\pm 2.00$
Fake-FDPET - $73.45\pm 3.01$ $61.05\pm 4.33$ $16.54\pm 3.20$ $4.07\pm 1.28$
LDPET-LDOS/ow - $71.09\pm 3.71$ $57.97\pm 3.15$ $17.39\pm 2.69$ $3.88\pm 1.09$
LDPET-LDOS - $73.97\pm 2.26$ $60.31\pm 2.72$ $14.00\pm 2.17$ $3.72\pm 0.46$
4.3 Comparisons with Previous Results
As shown in Table 3, Fig. 4 and 5, LDOS outperformed the state-of-the-art methods. Pre-training methods based on MAE [21, 16] did not yield significant improvements. LDOS uses semantic features extracted during the denoising process to achieve superior results. These findings are consistent with prior studies [21, 16, 18, 13], which indicate that pre-training on small-scale datasets provide limited advantages for medical imaging tasks.
4.4 Ablation Experiment
To evaluate the contribution of the proposed pipeline, we compared its against LDPET and fake FDPET. As shown in Table 4, LDOS improved Dice scores on both the 18F-FDG and 68Ga-FAPI datasets. Unlike fake FDPET, which employed sequential denoising and segmentation, LDOS’s collaborative design achieved superior results without complex denoising model training. LDOS reduces deployment costs while enhancing usability. Additionally, results from the LDPET-LDOS/ow LDPET-LDOS/ow (omits the L1 loss weight attenuation strategy) indicates that focusing on denoising semantic information during the early training stages and prioritizing segmentation optimization in the later stages positively impacts the model’s overall performance.
5 Conclusion
Ultra-LDPET organ segmentation is critical for reliable tracer uptake assessments and kinetic measurement, but remains largely unexplored. In this study, we propose LDOS, a CT-free ultra-LDPET organ segmentation pipeline. LDOS reinterprets LDPET as a naturally masked version of FDPET and extracts semantic information through a self-denoising process. By incorporating CT-derived organ annotations into the denoising process, LDOS improves anatomical boundary recognition and mitigates PET/CT misalignments, providing a novel solution for LDPET quantification analysis.